# TEA: Trajectory Encoding Augmentation for Robust and Transferable Policies in Offline Reinforcement Learning

### v1

@@ 1 Introduction & Related Work
Reinforcement Learning (RL) has seen notable success in addressing complex decision-making tasks across different fields. One ongoing challenge is to enable RL agents to generalize effectively when exposed to environments with varying dynamics, especially when the training needs to happen offline. This issue limits RL’s practical applications in real-world situations, where conditions may change unexpectedly. While transfer learning in RL usually seeks to tackle this issue by leveraging online data in both source and target environments, an open challenge remains: developing policies that are robust to new, unseen dynamics through adjusted offline training, purely with data from the source environments.
@@
Meta-reinforcement learning (Meta-RL) algorithms, such as Probabilistic Embeddings for Actor-Critic Reinforcement Learning (PEARL) [3] and Variational Bayes-Adaptive Deep RL (VariBAD) [4], learn latent task representations to facilitate rapid adaptation to new tasks through online interactions. However, such methods typically require additional environment interactions or task identifiers, which may not be feasible in offline settings.
@@ 2 Experimental Setup
This encoding is appended to original states, expanding the state space. To avoid additional data collection from the source environments, we retroactively extend states in the existing dataset by appending the relevant encodings.
@@ 3 Discussion
Results: The performance of each experiment on the new environment(s) after 20,000 timesteps of training on the source environments is shown in Table 1. In Figure 3 we see the performance of trajectory encoding augmentation (TEA) divided by the baseline performance for each individual new environment.

### v2

@@ 1 Introduction & Related Work
Reinforcement Learning (RL) has seen notable success in addressing complex decision-making tasks across different fields. One ongoing challenge is to enable RL agents to generalize effectively when exposed to environments with varying dynamics, especially when the training needs to happen offline. This issue limits RL’s practical applications in real-world situations, where conditions may change unexpectedly. While transfer learning in RL usually seeks to tackle this issue by leveraging online data in both source and target environments, an open challenge remains: developing policies to new, unseen dynamics through adjusted offline training, purely with data from the source environments.
@@
Meta-reinforcement learning (Meta-RL) algorithms, such as Probabilistic Embeddings for Actor-Critic Reinforcement Learning (PEARL) [3] and Variational Bayes-Adaptive Deep RL (VariBAD) [4], learn latent task representations to facilitate rapid adaptation to new tasks through online interactions. A Gaussian Process-based approach has been proposed to transfer knowledge across systems with slight variations by modeling correlations between members, focusing on multi-task settings [5]. In contrast, our method employs sequence encoders to capture latent dynamics, ensuring generalization in offline reinforcement learning without requiring task-specific identifiers.
@@ 2 Experimental Setup
This encoding is appended to the original states, expanding the state space. To avoid additional data collection from the source environments, we retroactively extend states in the existing dataset by appending the relevant encodings.
@@ 3 Discussion
Results: The performance of each experiment on the new environment(s) after 20,000 timesteps of training on the source environments is shown in Table 1. In Figure 3 we see the performance of trajectory encoding augmentation (TEA) divided by the baseline performance for each new environment.